OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. 3x - 2y = 9 and -3x + 2y = -9 Solve the system by using elimination or elimination with multiplication. One point for the steps and one point for the correct solution. 3x + 4y = -2 and 2x - 3y = 10

OpenStudy (anonymous):

3x + 4y = -2 2x - 3y = 10 3x + 4y = -2 x=(-2-4y)/3 2x - 3y = 10 2(-2-4y)/3 - 3y = 10 ( (-4-8y)/3-3y=10 ) * 3 -4-8y-9y=30 -4-17y=30 y=2

OpenStudy (anonymous):

3x - 2y = 9 and -3x + 2y = -9 u won't able to find the value of x and y with using the same equation

OpenStudy (anonymous):

you have infinite solutions to this equation ... all points lying on the line are solutions

OpenStudy (anonymous):

for the first one? u sure its not no solution?